# Latent heat estimation with machine learning

### v1

@@ Abstract
Predictions and phase probabilities. — The remaining 25% of the samples are used for predictions. Each configuration $i$ of size $L\times L$ sampled in MCPA simulations with energy $e$ is fed to the input of a trained network NN(L). The network outputs three numbers predicting membership of the ordered phase $p_{OS}^{i}(e)$, mixed phase $p_{MS}^{i}(e)$, and disordered phase $p_{DS}^{i}(e)$. The sum of these three numbers is equal to one. We repeat this process over the sample space, containing $N_{test}=2048$ samples with the given energy $e$. We calculated estimates of the probabilities $P_{xS}(e)$ averaging predictions $p_{xS}^{i}(e)$ over the sample space, $P_{xS}(e)=\sum_{i=1}^{N_{test}}p_{xS}^{i}(e)/N_{test}$, where $xS$ stands for one of the three phases, $OS,MS$, or $DS$. Thus, we obtain probability functions $P_{OS}(e)$, $P_{MS}(e)$, and $P_{DS}(e)$, with $P_{OS}(e)+P_{MS}(e)+P_{DS}(e)=1$.
@@
Figure 3: Probabilities of phases $P_{x}(E)$ for $L=60$ for 10-state Potts model, PM-10. Left panel is the training/testing with the raw dataset RD and right panel is the training/testing with the majority/minority dataset MD.
@@
At the same time, we found that the estimated probabilities contain some information about the details of the mixed phase. It is widely believed that there are four phase transitions in the mixed phase, which is a random mixture of ordered and disordered phases [14, 16]. In the case of the Potts model, these are blobs that reflect the ordered and disordered phases [27].
In the figure 5, the solid orange line corresponds to the mixed phase probability $P_{MS}$, the dotted green line corresponds to the disordered phase probability $P_{DS}$, and the dashed blue line corresponds to the ordered phase probability $P_{OS}$. The rightmost peak of $P_{OS}$ can be associated with the transition at energy $e_{1}(L)$ using the notations of Rose and Machta [16]. It is associated with the fluctuating droplets of OS phase within the DS sea, which vanishes at $e_{d}$. It is the precursor of the phase transition from disordered phase DS to mixed phase MS, as $e_{1}(L)\rightarrow e_{d}$ in the thermodynamic limit. By analogy, the leftmost peak of $P_{OD}$ can be associated with the transition at energy $e_{4}(L)$ and reflects DS droplets within the emerging sea of OS phase. Again, $e_{4}(L)\rightarrow e_{o}$ and the DS droplets completely vanishes at $e_{o}$. So, the transitions at $e_{1}(L)$ and $e_{4}(L)$ can be obtained only in the systems of finite size, and are not real phase transitions.

### v2

@@ Abstract
Predictions and phase probabilities. — The remaining 25% of the samples are used for predictions. Each configuration $i$ of size $L\times L$ sampled in MCPA simulations with energy $e$ is fed to the input of a trained network NN(L). The network outputs three numbers predicting membership of the ordered phase $p_{OS}^{i}(e)$, mixed phase $p_{MS}^{i}(e)$, and disordered phase $p_{DS}^{i}(e)$. The sum of these three numbers is equal to one. We repeat this process over the sample space, containing $N_{test}=2048$ samples with the given energy $e$. We calculated estimates of the probabilities $P_{xS}(e)$ averaging predictions $p_{xS}^{i}(e)$ over the sample space, $P_{xS}(e)=\sum_{i=1}^{N_{test}}p_{xS}^{i}(e)/N_{test}$, where $xS$ stands for one of the three phases, $OS,MS$, or $DS$. Thus, we obtain the probability functions $P_{OS}(e)$, $P_{MS}(e)$, and $P_{DS}(e)$, with $P_{OS}(e)+P_{MS}(e)+P_{DS}(e)=1$. Figures 3 and 4 show an example of phase probability estimation for model PM-10 and model PM-20, respectively, and present the results for both datasets, RD and MD. The network correctly predicts the phases of the models, and the smoothed sharp change in the probabilities near critical energies is explained by the finiteness of the system under study. The physics of some of the details is discussed in the last section “Discussion”.
@@
Figure 3: Probabilities of phases $P_{x}S(E)$ for $L=60$ for 10-state Potts model, PM-10. Left panel is the training/testing with the raw dataset RD and right panel is the training/testing with the majority/minority dataset MD.
@@
At the same time, we found that the estimated probabilities contain some information about the details of the mixed phase. It is widely believed that there are four phase transitions in the mixed phase, which is a random mixture of ordered and disordered phases [14, 16]. In the case of the Potts model, these are droplets that reflect the ordered and disordered phases [27].
In the figure 5, the solid orange line corresponds to the mixed phase probability $P_{MS}$, the dotted green line corresponds to the disordered phase probability $P_{DS}$, and the dashed blue line corresponds to the ordered phase probability $P_{OS}$. The rightmost peak of $P_{OS}$ can be associated with the transition at energy $e_{1}(L)$ using the notations of Rose and Machta [16]. It is associated with the fluctuating droplets of OS phase within the DS sea, which vanishes at $e_{d}$. It is the precursor of the phase transition from mixed phase MS to disordered phase DS, as $e_{1}(L)\rightarrow e_{d}$ in the thermodynamic limit. By analogy, the leftmost peak of $P_{OD}$ can be associated with the transition at energy $e_{4}(L)$ and reflects DS droplets within the emerging sea of OS phase. Again, $e_{4}(L)\rightarrow e_{o}$ and the DS droplets completely vanishes at $e_{o}$. So, the transitions at $e_{1}(L)$ and $e_{4}(L)$ can be obtained only in the systems of finite size, and are not real phase transitions.